OpenStudy (anonymous):

Which point lies on a circle that is centered at A(-3, 2) and passes through B(1, 3)? A.)C(-1, -2) B.)D(-6, 3) C.)E(-3, -3) D.)F(-2, 6)

OpenStudy (welshfella):

the general equation of a circle is (x - a)^2 + (y - b)^2 = r^2 where (a,b) is the center and r = radius

OpenStudy (welshfella):

so if the center is at (-3,2) we have (x + 3)^2 + (y - 2)^2 = r^2 Now we are given a points on the circle, (1,3) so if you plug these into the above you can find r^2

OpenStudy (welshfella):

can you do that?

OpenStudy (anonymous):

?